# Proper Effective Temperature and Order Parameters in Relativistic Non-Equilibrium Steady States

### v1

@@ V.1 Numerical analysis
In actual numerical calculations, appropriate cutoffs $\epsilon_{\text{UV}}$ and $\epsilon_{\text{IR}}$ are introduced for technical reasons: the boundary condition is specified at $u=u_{*}-\epsilon_{\text{IR}}$ and $m$ is read at $u=\epsilon_{\text{UV}}$, where $\epsilon_{\text{UV}}$ and $\epsilon_{\text{IR}}$ are appropriately chosen to sufficiently small values. We determine $m$ from $\theta^{\prime}(\epsilon_{\text{UV}})$, and the expectation value of the chiral condensate is obtained as $\theta^{\prime\prime}(\epsilon_{\text{UV}})/2$, as shown in (8).
V.2 Result
Fig. 1 shows the chiral condensate as a function of $v$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$.
@@ V.2 Result
The results of Fig. 1 suggest that a NESS at a higher value of $v$ mimics the chiral condensate of a higher value of $T$. Then let us check the dependence on $T_{\mathrm{eff}}$ of the chiral condensate.
Fig. 2 shows the chiral condensate as a function of $T_{\mathrm{eff}}$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$. Note that the value of $v$ is varying when we change $T_{\mathrm{eff}}$, since $T_{\mathrm{eff}}$ is a function of $v$.
@@
Fig. 3 shows the chiral condensate as a function of $T_{\mathrm{eff}}^{\mathrm{prop}}$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$. Note that the value of $v$ is varying when we change $T_{\mathrm{eff}}^{\mathrm{prop}}$, since $T_{\mathrm{eff}}^{\mathrm{prop}}$ is also a function of $v$.
@@
Fig. 4 and Fig. 5 show the chiral condensates as functions of $T_{\mathrm{eff}}^{\mathrm{prop}}$ at various values of $B$ and $\rho$, respectively.
@@
Now we have several comments. Naively, one may expect that natural parameters for the chiral condensate in NESS may be $(T,B,\rho,v)$ or $(T,B,\rho,T_{\mathrm{eff}})$. However, our results show only $(T_{\mathrm{eff}}^{\mathrm{prop}},B,\rho)$ are enough. When $v=0$, the system goes back to thermal equilibrium and $T_{\mathrm{eff}}^{\mathrm{prop}}$ agrees with the temperature of the heat bath $T$. Our observation that the chiral condensate in NESS can be described by $(T_{\mathrm{eff}}^{\mathrm{prop}},B,\rho)$ regardless of the value of $v$ means that the chiral condensate in NESS in the present setup can be predicted by using the behavior of the chiral condensate in equilibrium just by replacing $T$ with $T_{\mathrm{eff}}^{\mathrm{prop}}$  [^3] We have confirmed that $(T_{\mathrm{eff}}^{\mathrm{prop}},\rho)$ at the tricritical point, where the first-order phase transition turns into the second-order phase transition, at $B=1$ agrees with the results of Ref. [10] for equilibrium systems within numerical error when $T_{\mathrm{eff}}^{\mathrm{prop}}$ is identified with the heat-bath temperature of Ref. [10]. Note that the temperature $T$ in Ref. [10] is defined as $\pi$ times that in the present paper..
VI Conclusion and discussions
@@
We have found that the order parameter of the systems on the D5-brane is uniquely described by the proper effective temperature $T_{\mathrm{eff}}^{\mathrm{prop}}$, but not the naive effective temperature $T_{\mathrm{eff}}$. It is important to note that $T_{\mathrm{eff}}$ contains two elements. One is the effect of driving the system to a NESS, and the other is the relativistic effect of the system moving at a constant velocity $v$. This is because $T_{\mathrm{eff}}$ is the effective temperature measured by an observer on the rest frame of the heat bath. The transformation of $T_{\mathrm{eff}}$ into $T_{\mathrm{eff}}^{\mathrm{prop}}$ given by (28) removes the relativistic effect so that the net non-equilibrium effect is extracted [20]. $T_{\mathrm{eff}}^{\mathrm{prop}}$ is understood as the effective temperature measured by an observer on the rest frame of the system of interest on which NESS is realized. In this sense, $T_{\mathrm{eff}}^{\mathrm{prop}}$ is the NESS counterpart to the relativistic temperature introduced by Landsberg [21] and van Kampen [22], whereas $T_{\mathrm{eff}}$ is the NESS counterpart to the relativistic temperature proposed by Einstein [23] and Planck [24]. (See, e.g. Appendix A of [20].)
@@
We have several questions that should be clarified in future studies. One natural question is whether the properties of chiral condensate we have observed in the D3-D5 model in this study are general to other setups. Another question is whether a free energy can also be defined for the NESS studied here, and whether it can be described by $T_{\mathrm{eff}}^{\mathrm{prop}}$, $B$ and $\rho$  [^4] Discussions for the free energy of NESSs can be found, for example, in [33, 34]. . These issues are left for future research.
@@ Acknowledgements.
Appendix A Appendix A: Fluctuations in symmetric phase
We introduce a new coordinate $\xi$, defined by $d\xi^{2}/\xi^{2}=du^{2}/u^{2}(1-u^{4}/u_{h}^{4})$, so that the 6-dimensional part of the bulk geometry (1) can be written as
@@ Appendix A Appendix A: Fluctuations in symmetric phase
The new coordinates are related to the ones used in (2) via $l=\xi\cos\theta$ and $h=\xi\sin\theta$. Note that $d\xi^{2}=dl^{2}+dh^{2}$.
The D5-brane is allowed to fluctuate along the $(z,w^{4},w^{5},w^{6})$ directions. We define $\delta w$ as a fluctuation along the $(w^{4},w^{5},w^{6})$ directions. For instance, one may consider a fluctuation in $w^{4}$ while keeping $w^{5}=w^{6}=0$: we have three independent modes for $\delta w$. In the symmetric phase (i.e. when $\theta(u)=0$), the linearized equation of motion for $\delta w$ is given by (23), where

### v2

@@ 5.1 Numerical analysis
In actual numerical calculations, appropriate cutoffs $\epsilon_{\text{UV}}$ and $\epsilon_{\text{IR}}$ are introduced for technical reasons: the boundary condition is specified at $u=u_{*}-\epsilon_{\text{IR}}$ and $m$ is read at $u=\epsilon_{\text{UV}}$, where $\epsilon_{\text{UV}}$ and $\epsilon_{\text{IR}}$ are appropriately chosen to sufficiently small values. We determine $m$ from $\theta^{\prime}(\epsilon_{\text{UV}})$, and the expectation value of the chiral condensate is obtained as $\theta^{\prime\prime}(\epsilon_{\text{UV}})/2$, as shown in eq. (8).
5.2 Result
Figure 1 shows the chiral condensate as a function of $v$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$.
@@ 5.2 Result
The results of figure 1 suggest that a NESS at a higher value of $v$ mimics the chiral condensate of a higher value of $T$. Then let us check the dependence on $T_{\mathrm{eff}}$ of the chiral condensate.
Figure 2 shows the chiral condensate as a function of $T_{\mathrm{eff}}$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$. Note that the value of $v$ is varying when we change $T_{\mathrm{eff}}$, since $T_{\mathrm{eff}}$ is a function of $v$.
@@
Figure 3 shows the chiral condensate as a function of $T_{\mathrm{eff}}^{\mathrm{prop}}$ under $B=2$, $\rho=1$ for $T=0.02$, $0.06$, $0.10$, $0.14$, $0.18$ and $0.22$. Note that the value of $v$ is varying when we change $T_{\mathrm{eff}}^{\mathrm{prop}}$, since $T_{\mathrm{eff}}^{\mathrm{prop}}$ is also a function of $v$.
@@
Figure 4 and figure 5 show the chiral condensates as functions of $T_{\mathrm{eff}}^{\mathrm{prop}}$ at various values of $B$ and $\rho$, respectively.
@@
Now we have several comments. Naively, one may expect that natural parameters for the chiral condensate in NESS may be $(T,B,\rho,v)$ or $(T,B,\rho,T_{\mathrm{eff}})$. However, our results show only $(T_{\mathrm{eff}}^{\mathrm{prop}},B,\rho)$ are enough. When $v=0$, the system goes back to thermal equilibrium and $T_{\mathrm{eff}}^{\mathrm{prop}}$ agrees with the temperature of the heat bath $T$. Our observation that the chiral condensate in NESS can be described by $(T_{\mathrm{eff}}^{\mathrm{prop}},B,\rho)$ regardless of the value of $v$ means that the chiral condensate in NESS in the present setup can be predicted by using the behavior of the chiral condensate in equilibrium just by replacing $T$ with $T_{\mathrm{eff}}^{\mathrm{prop}}$. [^3] We have confirmed that $(T_{\mathrm{eff}}^{\mathrm{prop}},\rho)$ at the tricritical point, where the first-order phase transition turns into the second-order phase transition, at $B=1$ agrees with the results of ref. Evans:2011tk for equilibrium systems within numerical error when $T_{\mathrm{eff}}^{\mathrm{prop}}$ is identified with the heat-bath temperature of ref. Evans:2011tk . Note that the temperature $T$ in ref. Evans:2011tk is defined as $\pi$ times that in the present paper.
A natural question is whether the dependence of the order parameter on the proper effective temperature can be demonstrated analytically. Because the order parameter is extracted from $\theta(u)$ near the boundary, it would be sufficient to show that the equation of motion for $\theta$ depends explicitly on $T_{\mathrm{eff}}^{\mathrm{prop}}$ rather than on $T$ and $v$. Unfortunately, establishing this is far from trivial. To prove that equations of motion sharing the same proper effective temperature are equivalent, one must convert $T$ to $T_{\rm eff}^{\rm prop}$ through eqs. (24) and (26). However, eq. (24) contains $\theta(u_{*})$ and $\theta^{\prime}(u_{*})$ evaluated at $u_{*}$; hence the dependence of $T_{\rm eff}^{\rm prop}$ on $T$ enters through this $u_{*}$. On the other hand, $\theta(u_{*})$ and $\theta^{\prime}(u_{*})$ have to be selected so that $m$, which is given by the boundary value of $\theta(u)/u$, vanishes. Consequently, carrying out the conversion in advance would require knowing $\theta(u_{*})$ and $\theta^{\prime}(u_{*})$, information that only becomes available after the equation of motion has been solved. Our equation of motion is a nonlinear differential equation whose solution necessarily relies on numerical analysis. For this technical reason we have so far been unable to demonstrate analytically that two equations with different $T$ and $v$ but common $T_{\rm eff}^{\rm prop}$, $B$, and $\rho$ are indeed equivalent.
Figure 6: Comparison between the numerical solutions for two distinct parameter sets sharing the same proper effective temperature, $T_{\rm eff}^{\rm prop}=0.2052$. The brane embeddings are plotted in $(l,h)$ coordinates (see appendix A). The dotted lines mark the respective effective horizons. The blue dashed curve is the solution for a moving system with $(T,v)=(0.2,0.3)$ and $T_{\rm eff}^{\rm prop}=0.2052$. The red solid curve is for a static system ($v=0$) where the heat bath temperature $T$, which is also the proper effective temperature of the static system, is set to the $T_{\rm eff}^{\rm prop}$ of the moving system, i.e. $T=T_{\rm eff}^{\rm prop}=0.2052$. Both systems have $B=3$ and $\rho=0.02$. Despite the different effective horizon locations, the two brane profiles coincide to high accuracy, strongly suggesting that the system’s configuration is governed by $T_{\rm eff}^{\rm prop}$ rather than by $T$ and $v$ independently.
Numerically, however, once the proper effective temperature is matched, the solutions of the two equations completely overlap. Figure 6 compares representative numerical solutions of the two equations of motion for the parameter set (T,v)=(0.2, 0.3)(T,v)=(0.2,\,0.3) and (T,v)=(0.2052, 0)(T,v)=(0.2052,\,0) with common $(T_{\rm eff}^{\rm prop},B,\rho)=(0.2052,\,3,\,0.02)$. A noteworthy subtlety is that the effective geometries underlying the two formulations place the effective horizons at different radial positions. Despite this difference, the brane profiles outside the both horizons appear to coincide to high accuracy, suggesting that the two equations are analytically equivalent.
6 Conclusion and discussions
@@
We have found that the order parameter of the systems on the D5-brane is uniquely described by the proper effective temperature $T_{\mathrm{eff}}^{\mathrm{prop}}$, but not the naive effective temperature $T_{\mathrm{eff}}$. It is important to note that $T_{\mathrm{eff}}$ contains two elements. One is the effect of driving the system to a NESS, and the other is the relativistic effect of the system moving at a constant velocity $v$. This is because $T_{\mathrm{eff}}$ is the effective temperature measured by an observer on the rest frame of the heat bath. The transformation of $T_{\mathrm{eff}}$ into $T_{\mathrm{eff}}^{\mathrm{prop}}$ given by eq. (26) removes the relativistic effect so that the net non-equilibrium effect is extracted Hoshino:2018vne . $T_{\mathrm{eff}}^{\mathrm{prop}}$ is understood as the effective temperature measured by an observer on the rest frame of the system of interest on which NESS is realized. In this sense, $T_{\mathrm{eff}}^{\mathrm{prop}}$ is the NESS counterpart to the relativistic temperature introduced by Landsberg Landsberg_1966 and van Kampen vanKampen1968 , whereas $T_{\mathrm{eff}}$ is the NESS counterpart to the relativistic temperature proposed by Einstein einstein1907uber and Planck planck1907 . (See, e.g. appendix A of ref. Hoshino:2018vne .)
@@
We have several questions that should be clarified in future studies. A key question is whether we can show analytically that the order parameter or the equation of motion itself depends only on the proper effective temperature $T_{\rm eff}^{\rm prop}$ when $B$ and $\rho$ are fixed. Another question is whether the properties of chiral condensate we have observed in the D3-D5 model in this study are general to other setups. It is also important to see whether a free energy can also be defined for the NESS studied here, and whether it can be described by $T_{\mathrm{eff}}^{\mathrm{prop}}$, $B$ and $\rho$. [^4] Discussions for the free energy of NESSs can be found, for example, in ref. derrida2001free ; derrida2002large . These issues are left for future research.
@@ Acknowledgements.
Appendix A Fluctuations in symmetric phase
We introduce a new coordinate $\xi$, defined by $d\xi^{2}/\xi^{2}=du^{2}/u^{2}(1-u^{4}/u_{h}^{4})$, so that the 6-dimensional part of the bulk geometry eq. (1) can be written as
@@ Appendix A Fluctuations in symmetric phase
The new coordinates are related to the ones used in eq. (2) via $l=\xi\cos\theta$ and $h=\xi\sin\theta$. Note that $d\xi^{2}=dl^{2}+dh^{2}$.
The D5-brane is allowed to fluctuate along the $(z,w^{4},w^{5},w^{6})$ directions. We define $\delta w$ as a fluctuation along the $(w^{4},w^{5},w^{6})$ directions. For instance, one may consider a fluctuation in $w^{4}$ while keeping $w^{5}=w^{6}=0$: we have three independent modes for $\delta w$. In the symmetric phase (i.e. when $\theta(u)=0$), the linearized equation of motion for $\delta w$ is given by eq. (23), where